# Nyström Kernel Stein Discrepancy

### v1

@@ 4.3 Guarantees
The existing analysis of Nyström estimators [Rudi et al., 2015, Chatalic et al., 2022, Sterge and Sriperumbudur, 2022, Kalinke and Szabó, 2023] considers bounded kernels only. Indeed, if $\sup_{\mathbf{x}\in\mathbb{R}^{d}}\left\|{h_{p}\left(\cdot,{\mathbf{x}}\right)}\right\|_{\mathcal{H}_{h_{p}}}<\infty$, the consistency of (12) is implied by Chatalic et al. [2022, Theorem 4.1], which we include here for convenience of comparison.
@@
Assume the setting of Lemma 1, $C_{\mathbb{Q},h_{p}}\neq 0$, $m\geq 4$ Nyström samples, and bounded Stein feature map ($\sup_{\mathbf{x}\in\mathbb{R}^{d}}\left\|{h_{p}\left(\cdot,{\mathbf{x}}\right)}\right\|_{\mathcal{H}_{h_{p}}}=:K<\infty$). Then, for any $\delta\in(0,1)$ it holds with probability at least $1-\delta$ that
$\left|S_{p}(\mathbb{Q})-\tilde{S}_{p}\left(\hat{\mathbb{Q}}_{n}\right)\right|\leq\frac{c_{1}}{\sqrt{n}}+\frac{c_{2}}{m}+\frac{c_{3}\sqrt{\log(m/\delta)}}{m}\sqrt{\mathcal{N}_{\mathbb{Q},h_{p}}\left(\frac{12K^{2}\log(m/\delta)}{m}\right)},$ (16)
@@
Assume that the Stein feature map $h_{p}\left(\cdot,{X}\right)$ and the sampling distribution $\mathbb{Q}\in\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)$ are (i) sub-Gaussian in the sense of (2), that is, the condition $\left\|{\left\langle{h_{p}\left(\cdot,{X}\right),u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\lesssim\left\|{\left\langle{h_{p}\left(\cdot,{X}\right),u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}<\infty$ holds for all $u\in\mathcal{H}_{h_{p}}$, and (ii) it holds that $0<\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\lesssim\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}<\infty$.
@@
Let Assumption 1 hold, $C_{\mathbb{Q},h_{p}}\neq 0$, assume the setting of Lemma 1, $m\geq 4$ Nyström samples, and $\mathbb{P}=\mathbb{Q}$. Then, for any $\delta\in(0,1)$ it holds with probability at least $1-\delta$ that
$\left|S_{p}(\mathbb{Q})-\tilde{S}_{p}\left(\hat{\mathbb{Q}}_{n}\right)\right|\lesssim\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}\log(6/\delta)}{n}+\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)\log(6/\delta)}{n}}+$ (18)
$\quad+\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)\log(12n/\delta)\log(12/\delta)}}{m}\sqrt{\mathcal{N}_{\mathbb{Q},h_{p}}\left(\frac{c\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}{m}\right)}$ (19)
@@
$\left|S_{p}(\mathbb{Q})-\tilde{S}_{p}\left(\hat{\mathbb{Q}}_{n}\right)\right|=\mathcal{O}_{\mathbb{Q}}\!\left(\frac{1}{\sqrt{n}}\right),$ (20)
@@
    Sub-Gaussian assumption. Key to the proof of Theorem 2 is having an adequate notion of non-boundedness of the feature map. One approach—common for controlling unbounded real-valued random variables— is to impose a sub-Gaussian assumption. In Hilbert spaces, various definitions of sub-Gaussian behavior have been investigated [Talagrand, 1987, Fukuda, 1990, Antonini, 1997]; see Giorgobiani et al. [2020] for a recent survey. Classically, these require centered random variables. We note that with (6) the mean-zero requirement is satisfied for $\mathbb{P}=\mathbb{Q}$. Among the definitions of sub-Gaussianity, we carefully selected (i) Koltchinskii and Lounici [2017, Def. 2], and (ii) Assumption 1(ii). [^5] The former condition is also referred to as sub-Gaussian in Fukuda’s sense [Giorgobiani et al., 2020, Def. 1]. Specifically, (ii) is needed for our key Lemma B.3 which roughly states that sub-Gaussian vectors whitened by $C_{\mathbb{Q},h_{p},\lambda}^{-1/2}$ inherit the sub-Gaussian property, which opens the door for proving Theorem 2.
@@ 5 Experiments
We verify the viability of our proposed method, abbreviated as N-KSD in this section, by comparing its runtime and its power to existing methods: the quadratic time KSD [Liu et al., 2016, Chwialkowski et al., 2016], the linear-time goodness-of-fit test finite set Stein discrepancy (FSSD; Jitkrittum et al. 2017), and the near-linear-time goodness-of-fit test using random feature Stein discrepancy (L1 IMQ, L2 SechExp; Huggins and Mackey 2018). [^6] All the code replicating our experiments is available at https://github.com/FlopsKa/nystroem-ksd. For FSSD, we consider randomized test locations (FSSD-rand) and optimized test locations (FSSD-opt); the optimality is meant w.r.t. a power proxy detailed in the cited work [Jitkrittum et al., 2017]. For all competitors, we use the settings and implementations provided by the respective authors. We do not compare against RFF KSD, as Huggins and Mackey [2018] show that RFSDs, to which we compare against, achieve a better performance on the same set of experiments. We use the well-known Gaussian kernel $k(\mathbf{x},\mathbf{y})=\exp\left(-\gamma\left\|{\mathbf{x}-\mathbf{y}}\right\|_{\mathbb{R}^{d}}^{2}\right)$ ($\gamma>0$) with the median heuristic [Garreau et al., 2018], and the IMQ kernel $k(\mathbf{x},\mathbf{y})=\left(c^{2}+\left\|{\mathbf{x}-\mathbf{y}}\right\|_{\mathbb{R}^{d}}^{2}\right)^{\beta}$ [Gorham and Mackey, 2017], with the choices of $\beta<0$ and $c>0$ detailed in the respective experiment description. To approximate the null distribution of N-KSD, we perform a bootstrap with (15), setting $D=500$. To allow an easy comparison, our experiments replicate goodness-of-fit testing experiments from Chwialkowski et al. [2016], Jitkrittum et al. [2017] and Huggins and Mackey [2018]. We ran all experiments on a PC with Ubuntu 20.04, 124GB RAM, and 32 cores with 2GHz each.
Runtime. We set $m=\sqrt{n}$ for N-KSD. As per recommendation, we fix the number of test locations $J=10$ for L1 IMQ, L2 SechExp, and both FSSD methods. The runtime, see Figure 1(a) for the average over 10 repetitions (the error bars indicate the estimated $95\%$ quantile), behaves as predicted by the complexity analysis. The proposed approach runs orders of magnitudes faster than the quadratic time KSD estimator (7), and, for small samples sizes ($n\leq 500$), N-KSD is the fastest among all tested methods. From $n=2000$, all (near-)linear-time approaches are faster (excluding FSSD-opt, which has a relatively large fixed cost). Still, N-KSD achieves competitive runtime results even for $n=5000$.
@@
Restricted Boltzmann machine (RBM). Similar to Liu and Wang [2016], Jitkrittum et al. [2017], we consider the case where the target $\mathbb{P}$ is the non-normalized density of an RBM; the samples ^$\hat{\mathbb{Q}}_{n}$ are obtained from the same RBM perturbed by independent Gaussian noise with variance $\sigma^{2}$. For $\sigma^{2}=0$, $H_{0}:\mathbb{Q}=\mathbb{P}$ holds, and for $\sigma^{2}>0$, implying that the alternative $H_{1}:\mathbb{Q}\neq\mathbb{P}$ holds, the goal is to detect that the $n=1000$ samples come from a forged RBM. For the IMQ kernel (L1 IMQ, N-KSD IMQ, KSD IMQ), we set $c=1$ and $\beta=-1/2$. We show the results in Figure 1(d), using $100$ repetitions to obtain the power. KSD with the IMQ and with the Gaussian kernel performs best. Our proposed Nyström-based method ($m=4\sqrt{n}$) nearly matches its performance with both kernels while requiring only a fraction of the runtime. All other approaches achieve less power for $\sigma\in\{0.02,0.04\}$.
@@ A.2 Proof of Theorem 2
    By Assumption 1(ii), $Y$ is sub-Gaussian; hence it is also sub-exponential (Lemma D.2(3)), and therefore (Lemma B.2) it satisfies the Bernstein condition
@@
    for any $p\geq 2$. Notice that
    $B$ $=\left\|{Y}\right\|_{\psi_{1}}\stackrel{{\scriptstyle(a)}}{{\lesssim}}\left\|{Y}\right\|_{\psi_{2}}\stackrel{{\scriptstyle(b)}}{{\lesssim}}\left\|{Y}\right\|_{L_{2}(\mathbb{Q})}\stackrel{{\scriptstyle(c)}}{{=}}\left(\mathbb{E}_{X\sim\mathbb{Q}}\left\|h_{p}(\cdot,X)\right\|_{\mathcal{H}_{h_{p}}}^{2}\right)^{1/2}$ (A.7)
    =$\stackrel{{\scriptstyle(d)}}{{=}}\left[\mathbb{E}_{X\sim\mathbb{Q}}\operatorname{tr}\left(h_{p}\left(\cdot,{X}\right)\otimes h_{p}\left(\cdot,{X}\right)\right)\right]^{1/2}\stackrel{{\scriptstyle(e)}}{{=}}\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)},$ (A.8)
    (a) follows from Lemma D.2(3), (b) is implied by Assumption 1(ii), (c) comes from the definition of $Y$, in (d) we used Lemma D.1. Steinwart and Christmann [2008, (A.32)] allows commuting $\operatorname{tr}$ and integration, which, with (1), yields (e). As $\sigma\asymp B$, we also got that $\sigma\lesssim\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}$.
    Having obtained a bound on the moments, we can apply Bernstein’s inequality for separable Hilbert spaces (Yurinsky 1995; recalled in Theorem D.4) to the centered $\eta_{i}=h_{p}\left(\cdot,{\mathbf{x}_{i}}\right)\in\mathcal{H}_{h_{p}}$-s ($i\in[n]$), and obtain that for any $\delta\in(0,1)$ it holds with probability at least $31-\delta/$ 3 that
    $\left\|{\mu_{h_{p}}-\hat{\mu}_{h_{p}}}\right\|_{\mathcal{H}_{h_{p}}}=\left\|{\frac{1}{n}\sum_{i=1}^{n}\eta_{i}}\right\|_{\mathcal{H}_{h_{p}}}\lesssim\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}\log(6/\delta)}{n}+\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)\log(6/\delta)}{n}}.$ (A.10)
  • Term $t_{2}$. Assume that $0<\lambda\leq\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}$. Then, we can handle the second term with Lemma B.1 and obtain that for any $\delta\in(0,1)$ it holds with probability at least $31-\delta/$ 3 that
    $\left\|{\left(I-P_{\mathcal{H}_{h_{p},m}}\right)C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}\lesssim\sqrt{\lambda}$ (A.11)
@@
    Randomness in $\mathbf{x}_{i}$-s: Let $Z_{i}:=\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{\mathbf{x}_{i}}\right)}\right\|_{\mathcal{H}_{h_{p}}}$ ($i\in[n]$) with $(\mathbf{x}_{i})_{i=1}^{n}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{Q}$. By Assumption 1(ii) and Lemma B.3, the $Z_{i}$-s are sub-Gaussian random variable provided that $0<\lambda\leq\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}$. Hence, by Lemma B.6, with probability at least $61-\delta/$ 6, it holds that
@@
    We now bound $\left\|{Z_{1}}\right\|_{\psi_{2}}^{2}$:
    $\left\|{Z_{1}}\right\|_{\psi_{2}}^{2}\stackrel{{\scriptstyle(a)}}{{=}}\left\|{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{\mathbf{x}_{1}}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}^{2}\stackrel{{\scriptstyle(b)}}{{\lesssim}}\left\|{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{\mathbf{x}_{1}}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}^{2}\stackrel{{\scriptstyle(c)}}{{=}}\operatorname{tr}\left(C_{\mathbb{Q},h_{p},\lambda}^{-1}C_{\mathbb{Q},h_{p}}\right).$ (A.18)
    (a) follows from the definition of $Z_{1}$. (b) is implied by Assumption 1(ii) and Lemma B.3. (c) comes from Lemma B.4 with $A:=C_{\mathbb{Q},h_{p},\lambda}^{-1}$. We have shown that
@@
    $C$ =$=\left\{\left(\left(\mathbf{x}_{i}\right)_{i=1}^{n},\left(i_{j}\right)_{j=1}^{m}\right)\;:\;t_{3}\leq K\frac{\sqrt{2\log(12/\delta)}}{\sqrt{m}}\right\}\subseteq A.$ (A.22)
@@
To conclude, we use decomposition (A.5), and union bound (A.10), (A.11), and (A.25). Further, we observe that $\operatorname{tr}\left(C_{\mathbb{Q},h_{p},\lambda}^{-1}C_{\mathbb{Q},h_{p}}\right)=\mathcal{N}_{\mathbb{Q},h_{p}}(\lambda)$, and obtain that with probability at least $1-\delta$ it holds that
$\left|S_{p}(\mathbb{Q})-\tilde{S}_{p}\left(\hat{\mathbb{Q}}_{n}\right)\right|$ $\lesssim\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}\log(6/\delta)}{n}+\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)\log(6/\delta)}{n}}+$ (A.26)
$\quad+\sqrt{\frac{\lambda\mathcal{N}_{\mathbb{Q},h_{p}}(\lambda)\log(12n/\delta)\log(12/\delta)}{m}}$ (A.27)

### v2

@@ 4.3 Guarantees
The existing analysis of Nyström estimators [Rudi et al., 2015, Chatalic et al., 2022, Sterge and Sriperumbudur, 2022, Kalinke and Szabó, 2023] considers bounded kernels only. Indeed, if $\sup_{\mathbf{x}\in\mathbb{R}^{d}}\left\|{h_{p}\left(\cdot,{\mathbf{x}}\right)}\right\|_{\mathcal{H}_{h_{p}}}<\infty$, the consistency of (12) is implied by Chatalic et al. [2022, Theorem 4.1], which we include here for convenience of comparison. In the following, we denote the randomness in the choice of Nyström samples by $(i_{j})_{j=1}^{m}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\text{Unif}([n])=:\Lambda$, which means that $\tilde{\mathbf{x}}_{j}=\mathbf{x}_{i_{j}}$ with $j\in[m]$.
@@
Assume the setting of Lemma 1, $C_{\mathbb{Q},h_{p}}\neq 0$, $m\geq 4$ Nyström samples, and bounded Stein feature map ($\sup_{\mathbf{x}\in\mathbb{R}^{d}}\left\|{h_{p}\left(\cdot,{\mathbf{x}}\right)}\right\|_{\mathcal{H}_{h_{p}}}=:K<\infty$). Then, for any $\delta\in(0,1)$ it holds that
$\left(\mathbb{Q}^{n}\otimes\Lambda^{m}\right)\Bigg{(}\left|S_{p}(\mathbb{Q})-\tilde{S}_{p}\left(\hat{\mathbb{Q}}_{n}\right)\right|\leq\frac{c_{1}}{\sqrt{n}}+\frac{c_{2}}{m}+\frac{c_{3}\sqrt{\log\frac{m}{\delta}}}{m}\sqrt{\mathcal{N}_{\mathbb{Q},h_{p}}\!\left(\frac{12K^{2}\log\frac{m}{\delta}}{m}\right)}\Bigg{)}\geq 1-\delta,$ (16)
@@
We assume that the Stein feature map $h_{p}\left(\cdot,{X}\right)$ with the sampling distribution $\mathbb{Q}\in\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)$ is sub-Gaussian in the sense of (2), that is, $\left\|{\left\langle{h_{p}\left(\cdot,{X}\right),u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\lesssim\left\|{\left\langle{h_{p}\left(\cdot,{X}\right),u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}<\infty$ for all $u\in\mathcal{H}_{h_{p}}$.
@@
Let Assumption 1 hold, $C_{\mathbb{Q},h_{p}}\neq 0$, assume the setting of Lemma 1, $m\geq 4$ Nyström samples, and $\mathbb{P}=\mathbb{Q}$. Then, for any $\delta\in(0,1)$ it holds that
$\left(\mathbb{Q}^{n}\otimes\Lambda^{m}\right)\Bigg{(}$ $\left|S_{p}(\mathbb{Q})-\tilde{S}_{p}\left(\hat{\mathbb{Q}}_{n}\right)\right|\lesssim\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}\log(6/\delta)}{n}+\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)\log(6/\delta)}{n}}+$ (18)
$\quad+\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)\log(12n/\delta)\log(12/\delta)}}{m}\sqrt{\mathcal{N}_{\mathbb{Q},h_{p}}\left(\frac{c\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}{m}\right)}\Bigg{)}\geq 1-\delta$ (19)
@@
$\left|S_{p}(\mathbb{Q})-\tilde{S}_{p}\left(\hat{\mathbb{Q}}_{n}\right)\right|=\mathcal{O}_{\mathbb{Q}^{n}\otimes\Lambda^{m}}\!\left(\frac{1}{\sqrt{n}}\right),$ (20)
@@
    Sub-Gaussian assumption. Key to the proof of Theorem 2 is having an adequate notion of non-boundedness of the feature map. One approach—common for controlling unbounded real-valued random variables— is to impose a sub-Gaussian assumption. In Hilbert spaces, various definitions of sub-Gaussian behavior have been investigated [Talagrand, 1987, Fukuda, 1990, Antonini, 1997]; see Giorgobiani et al. [2020] for a recent survey. Classically, these require centered random variables. We note that with (6) the mean-zero requirement is satisfied for $\mathbb{P}=\mathbb{Q}$. Among the definitions of sub-Gaussianity, we carefully selected Koltchinskii and Lounici [2017, Def. 2]. [^5] The condition is also referred to as sub-Gaussian in Fukuda’s sense [Giorgobiani et al., 2020, Def. 1]. Specifically, the condition allows us to derive our key Lemma B.3, which intuitively states that norms of sub-Gaussian vectors whitened by $C_{\mathbb{Q},h_{p},\lambda}^{-1/2}$ inherit the sub-Gaussian property, opening the door for proving Theorem 2.
@@ 5 Experiments
We verify the viability of our proposed method, abbreviated as N-KSD in this section, by comparing its runtime and its power to existing methods: the quadratic time KSD [Liu et al., 2016, Chwialkowski et al., 2016], the linear-time goodness-of-fit test finite set Stein discrepancy (FSSD; Jitkrittum et al. 2017), RFF-based KSD approximations [Huggins and Mackey, 2018], and the linear-time goodness-of-fit test using random feature Stein discrepancy (L1 IMQ, L2 SechExp; Huggins and Mackey 2018). [^6] All the code replicating our experiments is available at https://github.com/FlopsKa/nystroem-ksd. For FSSD, we consider randomized test locations (FSSD-rand) and optimized test locations (FSSD-opt); the optimality is meant w.r.t. a power proxy detailed in the cited work [Jitkrittum et al., 2017]. For all competitors, we use the settings and implementations provided by the respective authors. We use the well-known Gaussian kernel $k(\mathbf{x},\mathbf{y})=\exp\left(-\gamma\left\|{\mathbf{x}-\mathbf{y}}\right\|_{\mathbb{R}^{d}}^{2}\right)$ ($\gamma>0$) with the median heuristic [Garreau et al., 2018], and the IMQ kernel $k(\mathbf{x},\mathbf{y})=\left(c^{2}+\left\|{\mathbf{x}-\mathbf{y}}\right\|_{\mathbb{R}^{d}}^{2}\right)^{\beta}$ [Gorham and Mackey, 2017], with the choices of $\beta<0$ and $c>0$ detailed in the respective experiment description. To approximate the null distribution of N-KSD, we perform a bootstrap with (15), setting $D=500$. To allow an easy comparison, our experiments replicate goodness-of-fit testing experiments from Chwialkowski et al. [2016], Jitkrittum et al. [2017] and Huggins and Mackey [2018]. We ran all experiments on a PC with Ubuntu 20.04, 124GB RAM, and 32 cores with 2GHz each.
Runtime. We set $m=4\sqrt{n}$ for N-KSD to match the settings in our other experiments. As per recommendation, we fix the number of test locations $J=10$ for L1 IMQ, L2 SechExp, and both FSSD methods. The data is randomly generated and has $d=10$ dimensions. The runtime, see Figure 1(a) for the average over 10 repetitions (the error bars indicate the estimated $95\%$ quantile), behaves as predicted by the complexity analysis. The proposed approach runs orders of magnitudes faster than the quadratic time KSD estimator (7). From $n=1500$, all (near-)linear-time approaches are faster (excluding FSSD-opt, which has a relatively large fixed cost). Still, N-KSD achieves competitive runtime results even for $n=5000$.
@@
Restricted Boltzmann machine (RBM). Similar to Liu and Wang [2016], Jitkrittum et al. [2017], we consider the case where the target $\mathbb{P}$ is the non-normalized density of an RBM; the samples ^$\hat{\mathbb{Q}}_{n}$ are obtained from the same RBM perturbed by independent Gaussian noise with variance $\sigma^{2}$. For $\sigma^{2}=0$, $H_{0}:\mathbb{Q}=\mathbb{P}$ holds, and for $\sigma^{2}>0$, implying that the alternative $H_{1}:\mathbb{Q}\neq\mathbb{P}$ holds, the goal is to detect that the $n=1000$ samples come from a forged RBM. For the IMQ kernel (L1 IMQ, N-KSD IMQ, KSD IMQ), we set $c=1$ and $\beta=-1/2$. We show the results in Figure 1(d), using $100$ repetitions to obtain the power. KSD with the IMQ and with the Gaussian kernel performs best. Our proposed Nyström-based method ($m=4\sqrt{n}$) nearly matches its performance with the IMQ kernel while requiring only a fraction of the runtime. Besides Cauchy RFF and Gauss RFF, all other approaches achieve less power for $\sigma\in\{0.02,0.04\}$.
@@ A.2 Proof of Theorem 2
    By Assumption 1 and Lemma B.3, $Y$ is sub-Gaussian; hence it is also sub-exponential (Lemma D.2(3)), and therefore (Lemma B.2) it satisfies the Bernstein condition
@@
    for any $p\geq 2$. Notice that $B=\left\|{Y}\right\|_{\psi_{1}}\stackrel{{\scriptstyle(a)}}{{\lesssim}}\left\|{Y}\right\|_{\psi_{2}}\stackrel{{\scriptstyle(b)}}{{\lesssim}}\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}$. (a) follows from Lemma D.2(3) and (b) is implied by Lemma B.3. As $\sigma\asymp B$, we also got that $\sigma\lesssim\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}$.
    Having obtained a bound on the moments, we can apply Bernstein’s inequality for separable Hilbert spaces (Yurinsky 1995; recalled in Theorem D.4) to the centered $\eta_{i}=h_{p}\left(\cdot,{\mathbf{x}_{i}}\right)\in\mathcal{H}_{h_{p}}$-s ($i\in[n]$), and obtain that for any $\delta\in(0,1)$ it holds that
    $\mathbb{Q}^{n}\Bigg{(}\underbrace{\left\|{\mu_{h_{p}}-\hat{\mu}_{h_{p}}}\right\|_{\mathcal{H}_{h_{p}}}}_{(=\left\|{\frac{1}{n}\sum_{i=1}^{n}\eta_{i}}\right\|_{\mathcal{H}_{h_{p}}})}\lesssim\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}\log(6/\delta)}{n}+\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)\log(6/\delta)}{n}}\Bigg{)}\geq 1-\delta/3.\quad\quad$ (A.7)
    Note that (A.7) also holds with the measure $\mathbb{Q}^{n}\otimes\Lambda^{m}$, since the event considered in (A.7) has no randomness w.r.t. $\Lambda^{m}$.
  • Term $t_{2}$. Assume that $0<\lambda\leq\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}$. Then, we can handle the second term with Lemma B.1 and obtain that for any $\delta\in(0,1)$ it holds that
    $\left(\mathbb{Q}^{n}\otimes\Lambda^{m}\right)\left(\left\|{\left(I-P_{\mathcal{H}_{h_{p},m}}\right)C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}\lesssim\sqrt{\lambda}\right)\geq 1-\delta/3$ (A.8)
@@
    Randomness in $\mathbf{x}_{i}$-s: Let $Z_{i}:=\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{\mathbf{x}_{i}}\right)}\right\|_{\mathcal{H}_{h_{p}}}$ ($i\in[n]$) with $(\mathbf{x}_{i})_{i=1}^{n}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{Q}$. By Assumption 1 and Lemma B.3, the $Z_{i}$-s are sub-Gaussian random variables. Hence, by Lemma B.5, with probability at least $61-\delta/$ 6, it holds that
@@
    By Lemma B.3, $\left\|{Z_{1}}\right\|_{\psi_{2}}^{2}\lesssim\operatorname{tr}\left(C_{\mathbb{Q},h_{p},\lambda}^{-1}C_{\mathbb{Q},h_{p}}\right)$. We have shown that
@@
    $C$ =$=\left\{\left(\left(\mathbf{x}_{i}\right)_{i=1}^{n},\left(i_{j}\right)_{j=1}^{m}\right)\;:\;t_{3}\leq K\frac{\sqrt{2\log(12/\delta)}}{\sqrt{m}},\left(\mathbf{x}_{i}\right)_{i=1}^{n}\in B\right\}\subseteq A.$ (A.18)
@@
To conclude, we use decomposition (A.5), and union bound (A.7), (A.8), and (A.21). Further, we observe that $\operatorname{tr}\left(C_{\mathbb{Q},h_{p},\lambda}^{-1}C_{\mathbb{Q},h_{p}}\right)=\mathcal{N}_{\mathbb{Q},h_{p}}(\lambda)$, and obtain that
$\left(\mathbb{Q}^{n}\otimes\Lambda^{m}\right)\Bigg{(}\left|S_{p}(\mathbb{Q})-\tilde{S}_{p}\left(\hat{\mathbb{Q}}_{n}\right)\right|$ $\lesssim\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}\log(6/\delta)}{n}+\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)\log(6/\delta)}{n}}+$ (A.22)
$\quad+\sqrt{\frac{\lambda\mathcal{N}_{\mathbb{Q},h_{p}}(\lambda)\log(12n/\delta)\log(12/\delta)}{m}}\Bigg{)}\geq 1-\delta$ (A.23)